The waiting time for a bus has a uniform distribution between 0 and 10 minutes. What is the probability that the waiting time for this bus is less than 4 minutes on a given day?
Answer #1

Solution :

Given that,

a = 0

b = 10

P(x < c) = (c - a) / (b - a)

P(x < 4) = (4 - 0) / (10 - 0) = 0.25

Probability = 0.4